OpenStudy (nthenic_oftime):

medal and Fan! Verify the identity. (please show steps and some explanation) cot θ ∙ sec θ = csc θ

Nnesha (nnesha):

what's the definition of sec tehta ?? and cot theta = what ?

Nnesha (nnesha):

for this question you just need to know definition of sin cos and cot

Nnesha (nnesha):

it's okay if you don't know :=) just let me know

Nnesha (nnesha):

anywz \[\sin \theta=\frac{ 1 }{ \csc \theta }~~~\cos \theta =\frac{ 1 }{ \sec \theta }~~~~\tan \theta=\frac{1 }{ \cot \theta}\] or you can say tan = sin /cos tan and cos are reciprocal of each otehr so if tan =sin/cos then cot =cos /sin for ur question rewrite cot and sec interms of sin/cos